OpenStudy (anonymous):

solve the equation -6p - 31= 3p - 12

OpenStudy (anonymous):

1) add \(6p\) to both side 2) add 12 to both sides 3) divide both sides by 9

OpenStudy (anonymous):

-6p-3p=31-12; -9p=19; p=-19/9